1. Gifblaar is a small South African shrub and one of the most poisonous plants knownbecauseit contains fluoroacetic acid (FCH2COOH), which has a pKaof 2.59.
Calculate the initial concentration of fluoroacetic acid in a solution that has the same pH as a 0.00135 M solution of HCl. answer in M
1
Expert's answer
2020-04-10T13:37:00-0400

HCl is a strong acid and dissociates completely to H+ and Cl-, therefore

pH = -lg[H+] = -lgC(HCl) = -lg0.00135 = 2.87

The equilibrium of dissociation is FCH2COOH = FCH2COO- + H+

the equilibrium constant is Ka = [H+]*[FCH2COO-]/[FCH2COOH]

C is initial concentration of the acid,

equilibrium concentrations are equal for [H+] = [FCH2COO-] = 10-2.87 = 0.00135 M,

equilibrium concentration of the acid is [FCH2COOH] = C - [H+] = C - 0.00135 M, hence:

Ka = [H+]2/(C-[H+])

10(-2.59) = 0.00257 = (0.00135)2/(C-0.00135),

C = (0.00135)2/0.00257 + 0.00135 = 0.00206 M

The initial concentration is 2.06*10-3 M